Solomon Corporation reported the following operating results for two consecutive years:

Required

Compute the percentage changes in Solomon Corporation’s income statement components between the two years. (Negative amounts should be indicated by a minus sign. Round your answers to 1 decimal place. (i.e., .234 should be entered as 23.4).)

SOLOMON CORPORATION
Income Statements
2018 2017 Percentage Change
Sales $1,295,000 $1,002,500 %
Cost of goods sold 796,425 601,500
Gross margin 498,575 401,000
Operating expenses 304,500 198,000
Income before taxes 194,075 203,000
Income taxes 61,400 52,800
Net income (loss) $132,675 $150,200 %

Zachary Company reported the following operating results for two consecutive years:

Required

Compute each income statement component for each of the two years as a percent of sales. (Percentages may not add exactly due to rounding. Round your answers to 1 decimal place. (i.e., .234 should be entered as 23.4).)

ZACHARY COMPANY
Vertical Analysis of Income Statements
2019 Percentage of Sales 2018 Percentage of Sales
Sales $1,002,000 % $1,077,000 %
Cost of goods sold 551,100 598,400
Gross margin on sales    450,900 478,600
Operating expenses 129,800 148,400
Income before taxes 321,100 330,200
Income taxes 80,500 80,600
Net income $240,600 % $249,600 %

Selected financial information for Rundle Company for 2019 follows

Sales $ 1,650,000
Cost of goods sold 1,155,000
Merchandise inventory
Beginning of year 160,000
End of year 197,000

Required

Assuming that the merchandise inventory buildup was relatively constant, how many times did the merchandise inventory turn over during 2019? (Round your answer to 2 decimal places.)

Merchandise inventory turnover times
Answer #1
1
Compute the percentage changes in Solomon Corporation’s income statement components between the two years. (Negative amounts should be indicated by a minus sign. Round your answers to 1 decimal place. (i.e., .234 should be entered as 23.4).
SOLOMON CORPORATION
Income Statements
2018 2017 Percentage Change
Sales $1,295,000 $1,002,500 29.2 %
Cost of goods sold 796,425 601,500 32.4 %
Gross margin 498,575 401,000 24.3 %
Operating expenses 304,500 198,000 53.8 %
Income before taxes 194,075 203,000 -4.4 %
Income taxes 61,400 52,800 16.3 %
Net income (loss) $132,675 $150,200 -11.7 %
Example: 100/1002500*1295000 = 29.2% sales increased in 2018 compared to 2017
Example: 100/601500*796425 = 32.4% cost of sales increased in 2018 compared to 2017
2
ZACHARY COMPANY
Vertical Analysis of Income Statements
2019 Percentage of Sales 2018 Percentage of Sales
Sales $1,002,000 100% $1,077,000 100%
Cost of goods sold 551,100 55% 598,400 56%
Gross margin on sales    450,900 45% 478,600 44%
Operating expenses 129,800 13% 148,400 14%
Income before taxes 321,100 32% 330,200 31%
Income taxes 80,500 8% 80,600 7%
Net income $240,600 24% $249,600 23%
Example: 100/1002500*551100= 55%
Example: 100/1077000*598400= 56%
3
Calculations:
Average inventory = (160,000+197,000)/2
178500
Inventory turnover
= Cost of goods sold / Average inventory
= 1,155,000 /178500 = 6.47